an iron rod is length 50m when the temperature is 20 degree celcuis by how much will it expand when the temperature rises to 30 degree celcuis
1
Expert's answer
2019-04-11T09:41:11-0400

By the definition of the linear thermal expansion we have:


"\Delta L = \alpha L_0 \Delta T,"

here, "\Delta L" is the difference in the length of the iron rod after the change in the temperature (or the expansion of the rod) , "L_0 = 50 m" is the original length of the iron rod, "\alpha = 11.8 \cdot 10^{-6} \dfrac{1}{^{\circ}C}" is the coefficient of linear expansion for the iron rod and "\Delta T = T_{final} - T_{initial} = 30^{\circ}C - 20^{\circ}C = 10^{\circ}C" is the change in temperature.

Then, from this formula we can calculate the linear expansion of the iron rod:


"\Delta L = 11.8 \cdot 10^{-6} \dfrac{1}{^{\circ}C} \cdot 50 m \cdot 10^{\circ}C = 5.9 \cdot 10^{-3} m."

Answer:

"\Delta L = 5.9 \cdot 10^{-3} m."